OpenStudy (anonymous):

If 10,800 cm2 of material is available to make a box with a square base and an open top, find the largest possible volume of the box.

OpenStudy (cwrw238):

surface area of box = x^2 + 4xh = 10800 whereby = side of base and h = height of box h = (10800 - x^2) / 4x V = x^2 * h V = x(10800 - x^2) / 4 V = 2700x - 0.25 x^2 now you can use calculus to find maximum volume

OpenStudy (cwrw238):

* where x = side of base

OpenStudy (cwrw238):

ooh - error in last line V = 2700x - 0.25x^3